OpenStudy (anonymous):

Solve the following system. x − y = 10 x + y = 4 (−3, 7) (−7, 3) (7, 3) (7, −3)

OpenStudy (anonymous):

(7, -3)

OpenStudy (anonymous):

add both equations x − y = 10 x + y = 4 ----------- 2x=14 x=14/2 x=7 now plug x value into one of the equation 7+y=4 y=4-7 y=-3 so x=7 and y=-3